OpenStudy (doriann):

Find the unknown percent 30 is ? % of 60

OpenStudy (evoker):

For a is b% of c rewrite as a/c=b/100 so for your case 30/60=?/100

OpenStudy (mathstudent55):

Percent is a ratio out of 100. An unknown percent means "what number out of 100." Your unknown is the percent, so your unknown is x/100. What you know is 30 out of 60. The question is 30 out of 60 is the same as what out of 100? Let x represent "what", the unknown, and write the equation \(\dfrac{30}{60} = \dfrac{x}{100} \) Solve for x.